OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

myininaya (myininaya):

@sirizarry78 Please help the user to see why that is. Involve the person asking.

OpenStudy (anonymous):

@myininaya what do you mean

myininaya (myininaya):

I was just asking the person who gave you the answer (the answer I deleted) to explain it.

myininaya (myininaya):

I guess they do not want to since they haven't replied. I will help. First get the absolute value part by itself. It is being multiplied by -3, so divide both sides by -3. Let me know what you get when doing this first step.

OpenStudy (anonymous):

\[\left| 2x+6 \right|=\frac{ -12 }{-3 }=4,2x+6=\pm 4\] solve the two values of x.